# Hallucinate at the Last in Long Response Generation: A Case Study on Long Document Summarization

### v1

@@ Abstract
Large Language Models (LLMs) have significantly advanced text generation capabilities, including tasks like summarization, often producing coherent and fluent outputs. However, faithfulness to source material remains a significant challenge due to the generation of hallucinations. While extensive research focuses on detecting and reducing these inaccuracies, less attention has been paid to the positional distribution of hallucination within generated text, particularly in long outputs. In this work, we investigate where hallucinations occur in LLM-based long response generation, using long document summarization as a key case study. Focusing on the challenging setting of long context-aware long response generation, we find a consistent and concerning phenomenon: hallucinations tend to concentrate disproportionately in the latter parts of the generated long response. To understand this bias, we explore potential contributing factors related to the dynamics of attention and decoding over long sequences. Furthermore, we investigate methods to mitigate this positional hallucination, aiming to improve faithfulness specifically in the concluding segments of long outputs.
@@
[…]
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have pushed the boundaries of human-like language generation, particularly in tasks like text summarization Chang et al. (2024b). With their enhanced scale and sophisticated training, LLMs now produce summaries exhibiting remarkable coherence and fluency, approaching human-level quality Roit et al. (2023); Song et al. (2025). This capability is transforming how we interact with large volumes of text, making information more accessible Achiam et al. (2023); Grattafiori et al. (2024); Team et al. (2024); GLM et al. (2024).
Figure 1: Comparison of the faithfulness scores of summaries generated around 50 words, 200 words, and 800 words in length for a 6K-token length Wikipedia context using GPT-4o mini.
@@
A persistent and critical challenge faced by LLMs is the phenomenon of hallucination, wherein the generated content is unfaithful to or unsupported by the input context Huang et al. (2025). While extensive prior work has focused on detecting and mitigating hallucinations, a significant limitation is the lack of research into the spatial distribution of factual errors within the generated sequence. Related research has explored positional biases in how models process input contexts, notably the Lost in the Middle phenomenon Liu et al. (2024b); Wan et al. (2025). However, understanding the distribution of errors within the generated output itself, particularly for long responses from long contexts, is equally crucial for effective diagnosis and mitigation.
Motivated by this critical research gap, our work provides the first dedicated investigation into the spatial distribution of hallucinations during long-response generation, especially in the challenging setting of long document summarization. Unlike prior work primarily addressing long input contexts or short outputs Tang et al. (2023); Zhang et al. (2024a); Bishop et al. (2024), we tackle long context-aware long response generation, which requires processing lengthy inputs and maintaining faithfulness over a significantly long output sequence. Specifically, we pose a core research question: How frequently do hallucinations occur in long-form generation tasks such as document summarization, and when they do, where in the output are they most likely to appear?
Our analysis reveals a surprising and concerning trend: hallucinations tend to concentrate towards the end of the generated text, a phenomenon we term "hallucinate at the last." As shown in Figure 1, faithfulness (See Section 3.1 for evaluation metric.) significantly decreases towards the end for long summaries (e.g., 800 words). Contrary to expectations of uniform distribution, this distinct positional bias highlights a critical vulnerability in LLMs when generating extended text. To systematically analyze this phenomenon and its implications, our research addresses three core questions:
@@
To answer these questions, we first empirically characterize the positional distribution of hallucinations (RQ1). Building on this, we explore potential causes of this phenomenon related to the generative process (RQ2). Finally, we propose and evaluate mitigation strategies targeting late-stage hallucinations (RQ3). Our findings highlight the necessity of considering the spatial dimension of errors and provide insights for developing more robust long-form generation systems.
@@ 2 Generating Summaries
This section describes the experimental setup used to generate the summaries analyzed in our study. We prompt LLMs to generate summaries from original documents, varying both the context length and the output length.
@@ 2.1 Input Context Lengths
To analyze the impact of input length, we evaluate model performance across a range of context lengths, varying from approximately 1K tokens up to 16K tokens. The specific context lengths used in our experiments range from 1K to 8K tokens, with increments of 1K.
@@ 2.2 Output Length Categories
  • Short Output: Summaries with a length between 100 and 200 words.
  • Long Output: Summaries with a length up to 30% of the input context length.
This distinction allows us to compare hallucination patterns in standard summary lengths versus significantly longer generated responses. By including both short and long output categories, we aim to systematically analyze if and how the positional bias of hallucinations manifests and potentially becomes more severe as the length of the generated summary increases. See Appendix A for more details.
@@ 2.3 Overall Faithfulness
To provide a general understanding of the faithfulness levels observed in the generated long summaries before conducting a detailed positional analysis, we first report the overall faithfulness scores.
@@
As shown in Table 1, we report the overall faithfulness of long summaries generated by Llama3.1-8B-Instruct Grattafiori et al. (2024) on the Wikipedia dataset, with context lengths ranging from 1K to 8K. We employ the FineSurE Song et al. (2024) framework with GPT-4 Achiam et al. (2023) to compute the overall faithfulness scores. FineSurE evaluates summaries by comparing each sentence with the source context and identifying specific error types when inconsistencies are detected.
As shown in the results, the overall performance is consistent with the results on the CNNDM Hermann et al. (2015) dataset reported in the original FineSurE paper, where the faithfulness of summaries generated by the Llama3-70B-Instruct model is reported to be 85.5%. Building on this, We perform a more fine-grained evaluation of summary faithfulness in Section 3, which includes a positional analysis to examine how factual consistency varies across different parts of the output.
@@ 3 Where Do Hallucinations Most Frequently Occur?
Figure 2: Comparison of faithfulness scores for short summaries generated by different models across increasing output lengths on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
@@ 3.1 Experimental Setup
We conduct experiments using several state-of-the-art Large Language Models to ensure our findings are not model-specific. The models evaluated include Llama3.1-8B-Instruct Grattafiori et al. (2024), Mistral-7B-Instruct-v0.3 Albert Q. Jiang et al. (2023), Qwen2.5-7B-Instruct Yang et al. (2024a), and GPT-4o mini Achiam et al. (2023).
Figure 3: Comparison of faithfulness scores for long summaries generated by different models across increasing output lengths on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
@@
To examine whether the phenomenon is domain-specific, we analyze the spatial distribution of faithfulness within generated summaries across various domains. In addition to the Wikipedia dataset used in the main experiment, we include the arXiv Cohan et al. (2018), PubMed Cohan et al. (2018), and GovReport Huang et al. (2021) datasets for further evaluation.
@@ 3.2 Evaluation Metric
We adopt an evaluation metric based on atomic facts, which has shown a significant effectiveness in evaluating summary factuality Min et al. (2023); Tang et al. (2024); Scirè et al. (2024); Jing et al. (2024); Wei et al. (2024); Yu et al. (2024).
@@
$A_{i}=\{a_{i1},a_{i2},...,a_{iNi}\}$ (1)
Using an NLI model, each filtered atomic fact is then compared with the source document $D=\{d_{1},...,d_{M}\}$ in a pair-wise manner. We measure only the entailment score, and each atomic fact is assigned the highest entailment score among its comparisons:
$\textrm{score}(a_{ij})=\max_{d_{m}\in{D}}\textrm{{Entail}}(a_{ij},d_{m})$ (2)
Finally, the average of the entailment scores for a set of filtered atomic facts in a sentence is used as the sentence’s faithfulness score:
@@
See Appendix B for more experimental details.
@@
In addition to the faithfulness score, we propose a simple yet effective metric for analyzing positional discrepancies in generated outputs. Specifically, each summary is divided into five equal length bins, and the average faithfulness score of the atomic facts within each bin is computed. We define sensitivity as the difference between the faithfulness score of the last bin and the average of the first four bins. A positive sensitivity value indicates a tendency to hallucinate at the last, whereas a negative value suggests otherwise. The larger the sensitivity, the more pronounced the hallucination at the end of the output.
@@ 3.3 Hallucinate at the Last
The latter part of the summary exhibits the lowest level of faithfulness across nearly all models, and across all contexts and summary lengths.
We present the positional faithfulness discrepancies of both short and long summaries generated by different models on the Wikipedia dataset, with respect to varying input context lengths, in Figure 2 and Figure 3, respectively. As shown in Figure 2, short summaries generated by the three models—excluding Mistral—consistently exhibit a decline in faithfulness towards the latter positions of the summary, regardless of the input context length, with the lowest faithfulness scores mostly observed at the end. Interestingly, we found that the positional discrepancies in generated summaries are not significantly affected by the input context length. Specifically, the decline in faithfulness at the end of summaries remains similar between the 1K and 8K context length settings.
However, the Hallucinate at the Last phenomenon becomes even more severe in long summaries, as shown in Figure 3. As the length of the generated summary increases, the faithfulness scores toward the end continue to decline, eventually dropping below 0.6 when 45 sentences are generated from an 8K context.
We report the sensitivity results for long summaries generated by the Llama3.1 model in Table 2, and those for the Mistral model in Table 3. As shown in the results, summaries generated by Llama3.1 tend to exhibit the lowest faithfulness scores in the final bin, resulting in high sensitivity. In contrast, the Mistral model often shows the lowest faithfulness scores in the initial bins, with sensitivity frequently falling below zero. We analyze the underlying causes of these contrasting patterns in Section 4. See Appendix C for more results.
@@
Llama3.1-8B-Instruct 1K 0.77 0.74 0.69 0.69 0.69 3.3
2K 0.74 0.72 0.70 0.66 0.63 7.5
3K 0.75 0.70 0.69 0.65 0.58 11.8
4K 0.74 0.71 0.68 0.69 0.64 6.5
5K 0.73 0.72 0.72 0.68 0.60 11.3
6K 0.74 0.72 0.66 0.65 0.61 8.3
7K 0.73 0.70 0.69 0.66 0.62 7.5
8K 0.76 0.70 0.72 0.69 0.61 10.8
@@
Table 2: Comparison of faithfulness scores and Sensitivity for long summaries generated by Llama3.1-8B-Instruct across output bins on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
Model Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Mistral-7B-Instruct-v0.3 1K 0.78 0.80 0.84 0.82 0.81 0.0
2K 0.75 0.80 0.84 0.82 0.81 -0.8
3K 0.81 0.80 0.82 0.84 0.83 -1.3
4K 0.79 0.81 0.83 0.81 0.81 0.0
5K 0.80 0.84 0.86 0.85 0.79 4.8
6K 0.82 0.87 0.87 0.87 0.86 -0.3
7K 0.82 0.84 0.82 0.83 0.85 -2.3
8K 0.87 0.83 0.87 0.89 0.81 5.5
Table 3: Comparison of faithfulness scores and Sensitivity for long summaries generated by Mistral-7B-Instruct-v0.3 across output bins on the Wikipedia dataset, which context lengths ranging from 1K to 8K.
@@
The latter part of the summary consistently demonstrates the lowest level of faithfulness across nearly all datasets, contexts, and summary lengths.
@@
We present the positional faithfulness discrepancy of long summaries generated by the Llama3.1-8B-Instruct model across different datasets in Figure 4. As the results indicate, the Hallucinate at the Last tendency is consistently observed across multiple domains. Faithfulness scores tend to decrease toward the final segments as the output length increases. In particular, as the length of the generated summary increases, the faithfulness of its later parts continues to decline. Notably, for the Pubmed dataset, the faithfulness score falls below 0.5 when the model generates 45 sentences at a context length of 8K. Additional sensitivity results across diverse datasets are reported in Appendix D. Notably, across all datasets, none of the computed sensitivity scores fall below zero.
@@
Figure 4: Comparison of faithfulness scores in long summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths in multiple domains, with context lengths ranging from 1K to 8K. We report results in short summaries in Appendix D.

### v2

@@ Abstract
Large Language Models (LLMs) have significantly advanced text generation capabilities, including tasks like summarization, often producing coherent and fluent outputs. However, faithfulness to the source material remains a significant challenge due to the generation of hallucinations by LLMs. While extensive research focuses on detecting and reducing these inaccuracies, less attention has been paid to the positional distribution of hallucinations within generated text, particularly in long outputs. In this work, we investigate where hallucinations occur in LLM-based long response generation, using long document summarization as a key case study. Focusing on the challenging setting of long context-aware long response generation, we find a consistent and concerning phenomenon: hallucinations tend to concentrate disproportionately in the latter parts of the generated long response. To understand this bias, we explore potential contributing factors related to the dynamics of attention and decoding over long sequences. Furthermore, we investigate methods to mitigate this positional hallucination, aiming to improve faithfulness specifically in the concluding segments of long outputs.
@@
Joonho Yang1, Seunghyun Yoon2, Hwan Chang1, Byeongjeong Kim1, Hwanhee Lee1thanks: Corresponding author. 1Department of Artificial Intelligence, Chung-Ang University, 2Adobe Research, USA {plm3332, hwanchang, michael97k, hwanheelee}@cau.ac.kr, syoon@adobe.com
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have pushed the boundaries of human-like language generation, particularly in tasks like text summarization Chang et al. (2024b). With their enhanced scale and sophisticated training, LLMs now produce summaries exhibiting remarkable coherence and fluency, approaching human-level quality Song et al. (2025). This capability is transforming how we interact with large volumes of text, making information more accessible Achiam et al. (2023); Team et al. (2024).
Figure 1: Comparison of the faithfulness scores of summaries generated around 50, 200, and 800 words in length for a 6K-token length Wikipedia context using GPT-4o mini.
@@
A persistent and critical challenge faced by LLMs is the phenomenon of hallucination, wherein the generated content is unfaithful to or unsupported by the input context Huang et al. (2025). While extensive prior work has focused on detecting and mitigating hallucinations, a significant limitation is the lack of research into the positional distribution of factual errors within the generated sequence. Related research has explored positional biases in how models process input contexts, notably the Lost in the Middle phenomenon Liu et al. (2024b); Wan et al. (2025). However, understanding the distribution of errors within the generated output itself, particularly for long responses from long contexts, is equally crucial for effective diagnosis and mitigation.
Motivated by this critical research gap, our work provides the first dedicated investigation into the positional distribution of hallucinations during long-response generation, especially in the challenging setting of long document summarization. Unlike prior work primarily addressing long input contexts or short outputs Zhang et al. (2024a); Bishop et al. (2024), we tackle long context-aware long response generation, which requires processing lengthy inputs and maintaining faithfulness over a significantly long output sequence. Specifically, we pose a core research question: How frequently do hallucinations occur in long-form generation tasks such as document summarization, and when they do, where in the output are they most likely to appear?
Our analysis reveals a surprising and concerning trend: hallucinations tend to concentrate towards the end of the generated text, a phenomenon we term "hallucinate at the last." As shown in Figure 1, faithfulness significantly decreases towards the end for long summaries (e.g., 800 words). Contrary to expectations of uniform distribution, this distinct positional bias highlights a critical vulnerability in LLMs when generating extended text. To systematically analyze this phenomenon and its implications, our research addresses three core questions:
@@
To answer these questions, we empirically characterize the positional distribution of hallucinations (RQ1), investigate their generative causes (RQ2), and propose mitigation strategies (RQ3). Our results underscore the importance of the positional dimension of errors in building robust long-form generation systems.
@@ 2 Generating Summaries
This section describes the experimental setup used to generate the summaries analyzed in our study. We prompt LLMs to generate summaries from original documents, varying both the context length and the output length. For all experiments, we generate summaries using greedy decoding. The prompt is provided in Appendix A.
@@ 2.1 Input Context Lengths
To analyze the impact of input length, we evaluate model performance across a range of context lengths, varying from approximately 1K tokens up to 8K tokens. For the Llama model specifically, we further extend the evaluation to context lengths of 12K and 16K tokens.
@@ 2.2 Output Length Categories
  • Standard Summary: Summaries with a length between 100 and 200 words.  [^1] The length of the standard summaries in our study was determined based on the average length of the reference summaries in the dataset, which is employed in Section 4.
  • Long Summary: Summaries with a length up to 30% of the input context length.
This distinction allows us to compare hallucination patterns in standard summary lengths versus significantly longer generated responses. By including both standard and long output categories, we aim to systematically analyze if and how the positional bias of hallucinations manifests and potentially becomes more severe as the length of the generated summary increases. See Appendix A for more details.
@@ 2.3 Overall Faithfulness
As a preliminary step, we measure the overall faithfulness of the generated long summaries with FineSurE Song et al. (2024) to ensure consistent quality across different context lengths, prior to our main positional analysis. Table 1 shows results for Llama3.1-8B-Instruct Grattafiori et al. (2024) on the Wikipedia dataset (1K–8K context lengths), evaluated using FineSurE with GPT-4 Achiam et al. (2023). FineSurE assesses each sentence against the source, identifying error types when inconsistencies occur. The results align with those on CNNDM Hermann et al. (2015), where Llama3-70B-Instruct achieved 85.5% faithfulness. We next conduct a finer-grained positional analysis in Section 3 for various setups.
@@ 3 Where Do Hallucinations Most Frequently Occur?
Figure 2: Comparison of faithfulness scores for standard summaries generated by different models across increasing output lengths on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
@@ 3.1 Experimental Setup
We evaluate our approach on multiple state-of-the-art LLMs, including Llama3.1-8B-Instruct, Gemma3-12B-Instruct Team et al. (2025), Qwen2.5-7B-Instruct Yang et al. (2024a), and GPT-4o mini, to ensure model-agnostic findings. For Gemma, we retain the default sliding window attention context length of 1024 tokens, while for Qwen, we modify the default setting of over 131K tokens to 16K tokens.
@@
To examine whether the phenomenon is domain-specific, we analyze the positional distribution of faithfulness within generated summaries across various domains. In addition to the Wikipedia dataset used in the main experiment, we include the arXiv Cohan et al. (2018), PubMed Cohan et al. (2018), and GovReport Huang et al. (2021) datasets for further evaluation.
@@ 3.2 Evaluation Metric
To enable an in-depth analysis of positional bias within each generated summary, we use the MiniCheck-Flan-T5-Large Tang et al. (2024) model to evaluate faithfulness. MiniCheck is particularly suitable for this granular task, as it is an efficient and high-performing fact-checking model that leverages atomic facts Min et al. (2023); Jing et al. (2024); Yu et al. (2024).
@@
$A_{i}=\{a_{i1},a_{i2},...,a_{iN_{i}}\}$ (1)
Using a MiniCheck model, which is denoted as $\mathcal{M}$, each filtered atomic fact is then compared with the source document $D=\{d_{1},...,d_{M}\}$ in a pair-wise manner. Each atomic fact is assigned the highest score among its comparisons:
$\textrm{score}(a_{ij})=\max_{d_{m}\in{D}}\mathcal{M}(d_{m},a_{ij})$ (2)
Finally, we use the average of the scores for a set of filtered atomic facts in a sentence as its faithfulness score:
@@
Additional experimental details and an example of the evaluation procedure are provided in Appendix B.
Furthermore, we verify the quality of atomic facts generated by the LLM via human evaluation in Table 2 for atomic facts extracted from long summaries generated by Llama.
Context Length Validation(%)
1K 98.55
2K 99.08
3K 97.62
4K 95.92
5K 99.01
6K 96.18
7K 99.12
8K 99.64
Table 2: Human evaluation of the quality (completeness) of atomic facts. For each context length, we randomly sampled 10 summaries and conducted the evaluation manually. Each atomic fact was annotated as either true or false depending on whether it is present in the summary.
@@
In addition to the faithfulness score, we propose a simple yet effective metric for analyzing positional discrepancies in generated outputs. Specifically, following Singh and Goshtasbpour (2022), each summary is divided into five equal length bins or segments, and the average faithfulness score of the atomic facts within each bin is computed. We define sensitivity as the difference between the faithfulness score of the last bin and the average of the first four bins. A positive sensitivity value indicates a tendency to hallucinate at the last, whereas a negative value suggests otherwise. The larger the sensitivity, the more pronounced the hallucination at the end of the output.
Models Context Generated Summary Bins Sensitivity
Length 1 2 3 4 5
Llama3.1-8B-Instruct Overall Avg. 0.81 0.82 0.79 0.80 0.75 5.5
12K 0.80 0.79 0.66 0.63 0.69 9.5
16K 0.86 0.83 0.85 0.83 0.79 5.25
Llama3.1-70B-Instruct Overall Avg. 0.83 0.82 0.83 0.81 0.77 5.25
Gemma3-4B-Instruct Overall Avg. 0.85 0.82 0.83 0.85 0.79 4.75
Gemma3-12B-Instruct Overall Avg. 0.85 0.83 0.84 0.83 0.79 4.75
GPT-4o mini Overall Avg. 0.83 0.81 0.81 0.78 0.76 4.75
Qwen2.5-7B-Instruct Overall Avg. 0.86 0.83 0.83 0.82 0.84 -0.5
Table 3: Faithfulness scores and sensitivity for long summaries generated by various models on the Wikipedia dataset. Overall Avg. denotes the average faithfulness score aggregated across context lengths from 1K to 8K.
@@ 3.3 Hallucinate at the Last
Llama3.1-8B-Instruct 1K 0.92 0.90 0.91 0.87 0.68 22.00
2K 0.82 0.74 0.83 0.72 0.71 6.75
3K 0.94 0.94 0.81 0.86 0.68 20.75
4K 0.92 0.92 0.88 0.87 0.82 7.75
5K 0.91 0.88 0.89 0.78 0.70 16.50
6K 0.86 0.84 0.79 0.89 0.73 11.50
7K 0.97 0.89 0.82 0.94 0.66 24.50
8K 0.93 0.90 0.86 0.79 0.65 22.00
Table 4: Human evaluation of faithfulness scores and sensitivity for long summaries generated by Llama on the Wikipedia dataset. Each atomic fact was annotated as either true or false depending on whether it is present in the document. The results were manually evaluated and aggregated.
The latter part of the summary exhibits the lowest level of faithfulness across multiple models, consistently observed across various contexts and summary lengths. The decline becomes most pronounced when the context length reaches 8K and the summaries are long.
We present the positional faithfulness discrepancies of both standard and long summaries generated by different models on the Wikipedia dataset, with respect to varying input context lengths, in Figure 2 and Figure 3, respectively. As shown in Figure 2, standard summaries generated by the three models, excluding Qwen, consistently exhibit a decline in faithfulness towards the latter positions of the summary, regardless of the input context length, with the lowest faithfulness scores mostly observed at the end. Interestingly, we found that the positional discrepancies in generated summaries are not significantly affected by the input context length. Specifically, the decline in faithfulness at the end of summaries remains similar between the 1K and 8K context length settings.
@@
Figure 3: Comparison of faithfulness scores for long summaries generated by different models across increasing output lengths on the Wikipedia dataset, with context lengths ranging from 1K to 8K.
However, the Hallucinate at the Last phenomenon becomes even more severe in long summaries except for Qwen, as shown in Figure 3. For these models, as the length of the generated summary increases, the faithfulness scores toward the end continue to decline, eventually dropping below a faithfulness score of 0.70 when 45 sentences are generated from an 8K context. Qwen, in contrast, often exhibits the largest drop in faithfulness in the middle of the summary, after which the scores recover toward the end.
Tables 3 report the faithfulness and sensitivity of various models. We aggregate the faithfulness scores across context lengths from 1K to 8K and report the overall average. All models except Qwen show markedly low faithfulness scores in the final bin, reflecting high sensitivity. Consistent with the human evaluation in Table 4, Llama demonstrates the highest sensitivity regardless of model size or context length. Qwen, however, is the only model that records sensitivity values below zero. Unlike the consistent decline observed in the other models, Qwen maintains relatively stable faithfulness across bins, resulting in negative sensitivity, and we provide a detailed analysis of these contrasting patterns in Section 4. Supplementary results across context lengths and methods are provided in Appendix C.
@@
The latter part of the summary consistently demonstrates the lowest level of faithfulness across nearly all datasets, with a marked decline once the context length exceeds 7K.
@@
We present the positional faithfulness discrepancy of long summaries generated by the Llama model across different datasets in Figure 4. As the results indicate, the Hallucinate at the Last tendency is consistently observed across multiple domains. Faithfulness scores tend to decrease toward the final segments as the output length increases. In particular, as the length of the generated summary increases, the faithfulness of its later parts continues to decline, with the deterioration becoming especially pronounced once the context length exceeds 7K. Additional sensitivity results across diverse datasets are provided in Appendix D.
@@
Figure 4: Comparison of faithfulness scores in long summaries generated by the Llama3.1-8B-Instruct model across increasing output lengths in multiple domains, with context lengths ranging from 1K to 8K.